OpenStudy (anonymous):

Find The Missing Number 5x+4=24

OpenStudy (anonymous):

5x+4=24 24 - 4 = 20 20 / 5 = 4 x = 4

OpenStudy (anonymous):

Thats Good . Thank You

OpenStudy (anonymous):

your welcome + thanks you, as well.